# Align-SLM: Textless Spoken Language Models with Reinforcement Learning from AI Feedback

### v1

@@ 3.2.1 Continuation Likelihood: Perplexity
For the PPL of $N$ continuations ($PPL_{N}$), we first filter out the auto-BLEU $a_{i}$ higher than $\delta$, then select the lowest PPL sample as $y_{c}$. The threshold of the auto-BLEU score is selected by the score distribution between ground truth continuation and the generated result (Please see the Appendix G for details). The $y_{r}$ is the continuation with the highest PPL. The $(y_{c},y_{r})$ is created as below:
@@ 3.2.2 LLM Evaluation: Mistral Score
Instruction-tuned LLMs can follow instructions and understand semantics well (Chung et al., 2024). We propose using an LLM to judge the quality of the speech continuation, which evaluates the entire input and predicts the score. The prompt (see the Appendix for more details) is utilized to instruct the model to provide a score between 1 to 5 (1 denoting bad and 5 denoting good) based on the likelihood and meaningfulness of continuation given text prompt. Since we use the Mistral model for LLM evaluation, we call this “Mistral score", denoted as $s$. To let the model learn to distinguish the preferred and unpreferred continuations, a certain threshold is set for the Mistral score to ensure the difference in quality. $s_{c}$ is the threshold of the chosen sample, and $s_{r}$ is the threshold of the rejected sample. $s_{c}$ should be larger than $s_{r}$. auto-BLEU threshold is also used to recognize the naively repeated samples as rejected. We select the $(y_{c},y_{r})$ as below:
@@
The $s_{c}$ and $s_{r}$ values are selected based on a preliminary analysis of the SLM’s score distribution. For more details on the distribution of Mistral scores, please see Appendix H.
@@ 4.2.3 Generative Speech Continuation
GPT4-o score: GPT4-o (OpenAI, 2023) has shown remarkable text understanding performance and can serve as alternative human evaluators (Chiang and Lee, 2023a; Liu et al., 2023; Chiang and Lee, 2023b), showing high correlation with human judgments. We leverage GPT4-o as a proxy for human evaluations. Following llm evaluation (Chiang and Lee, 2023a), the instruction first analyzes the sentence and then provides the score from 1 to 5, to judge the semantic coherence, meaningfulness, and grammatical correctness of the ASR transcribed continuation given a prompt (1 denoting bad and 5 denoting good). The instruction prompt is shown in Appendix F.
@@ 5.3 Improvement of Curriculum Learning
After training the pre-trained model with the first iteration of Align-SLM, we consider the resulting model a stronger starting point and generate new preference data with more stringent preference data selection criteria. Results in Table 2 indicate that curriculum learning improves most metrics on the Zerospeech and StoryCloze benchmarks (rows 8 to 9 for 1.3B, rows 16 to 17 for 7B), particularly for T-StoryCloze, which requires fine-grained relevance between the speech story prompt and its continuation. Additionally, we observe improvements in speech continuation; for example, the GPT4-o score increases from 2.06 to 2.29 for the 1.3B model. We also experimented with further increasing the number of curriculum learning iterations, which continued to enhance performance (see the discussion in Appendix D).
@@ Appendix B Details of Subjective Evaluation
Appendix C Speech Continuation Examples
@@
Appendix D Number of Iteration for Curriculum Learning
@@
Appendix E Instruction for the Mistral LLM for evaluation
@@
Appendix F Instruction for the GPT4-o for evaluation
@@
Appendix G auto-BLEU Score Distribution
@@
Appendix H Mistral Score Distribution
@@
Appendix I Correlation Between GPT4-o score and MMOS

### v2

@@ 3.2.1 Continuation Likelihood: Perplexity
For the PPL of $N$ continuations ($PPL_{N}$), we first filter out the auto-BLEU $a_{i}$ higher than $\delta$, then select the lowest PPL sample as $y_{c}$. The threshold of the auto-BLEU score is selected by the score distribution between ground truth continuation and the generated result (Please see the Appendix H for details). The $y_{r}$ is the continuation with the highest PPL. The $(y_{c},y_{r})$ is created as below:
@@ 3.2.2 LLM Evaluation: Mistral Score
Instruction-tuned LLMs can follow instructions and understand semantics well (Chung et al., 2024). We propose using an LLM to judge the quality of the speech continuation, which evaluates the entire input and predicts the score. The prompt (see the Appendix for more details) is utilized to instruct the model to provide a score between 1 to 5 (1 denoting bad and 5 denoting good) based on the likelihood and meaningfulness of continuation given text prompt. Since we use the Mistral model for LLM evaluation, we call this “Mistral score", denoted as $s$. To let the model learn to distinguish the preferred and unpreferred continuations, a certain threshold is set for the Mistral score to ensure the difference in quality. $s_{c}$ is the threshold of the chosen sample, and $s_{r}$ is the threshold of the rejected sample. $s_{c}$ should be larger than $s_{r}$. The auto-BLEU threshold is also used to recognize the naively repeated samples as rejected. We select the $(y_{c},y_{r})$ as below:
@@
The $s_{c}$ and $s_{r}$ values are selected based on a preliminary analysis of the SLM’s score distribution. For more details on the distribution of Mistral scores, please see Appendix I. We provide an example to illustrate the preference data selection in the Appendix C section.
@@ 4.2.3 Generative Speech Continuation
GPT4-o score: GPT4-o (OpenAI, 2023) has shown remarkable text understanding performance and can serve as alternative human evaluators (Chiang and Lee, 2023a; Liu et al., 2023; Chiang and Lee, 2023b), showing high correlation with human judgments. We leverage GPT4-o as a proxy for human evaluations. Following llm evaluation (Chiang and Lee, 2023a), the instruction first analyzes the sentence and then provides the score from 1 to 5, to judge the semantic coherence, meaningfulness, and grammatical correctness of the ASR transcribed continuation given a prompt (1 denoting bad and 5 denoting good). The instruction prompt is shown in Appendix G.
@@ 5.3 Improvement of Curriculum Learning
After training the pre-trained model with the first iteration of Align-SLM, we consider the resulting model a stronger starting point and generate new preference data with more stringent preference data selection criteria. Results in Table 2 indicate that curriculum learning improves most metrics on the Zerospeech and StoryCloze benchmarks (rows 8 to 9 for 1.3B, rows 16 to 17 for 7B), particularly for T-StoryCloze, which requires fine-grained relevance between the speech story prompt and its continuation. Additionally, we observe improvements in speech continuation; for example, the GPT4-o score increases from 2.06 to 2.29 for the 1.3B model. We also experimented with further increasing the number of curriculum learning iterations, which continued to enhance performance (see the discussion in Appendix E).
@@ Appendix B Details of Subjective Evaluation
Appendix C Preference-based selection illustration
Given the speech prompt “a man was looking in from the corridor behind", we draw five candidate continuations from the model and assign each a Mistral score (higher is better). Table 4 lists the continuations and their scores. We keep a continuation if $s_{i}\geq s_{c}=3$, reject it if $s_{i}\leq s_{r}=1$, and otherwise mark it as filtered (neither accepted nor explicitly rejected). In this example, continuations (1) and (5) are kept; continuations (2) and (4) are rejected due to clear irrelevance; and continuation (3) is filtered. Notably, the lowest-scoring continuations (score 1) tend to be semantically unrelated to the prompt, underscoring the value of the LLM feedback.
Prompt: a man was looking in from the corridor behind
# Sampled speech continuation Score
1 He was seen, he stopped before the man. He said, I have a book you read. 3
2 When they drew near and received their letter, and looking down, showed them something of the startling effect of the diamond. 1
3 But every time he reached the open window he saw the little man pressing in a silk sash. 2
4 When he heard the dog talking to him. Said, My father, did you hear the dog? Said he, did you… 1
5 And the door opened, and he heard the words. And there they all began, Boyce said to himself. 3
Table 4: Prompt, sampled continuations, and their Mistral scores.
Appendix D Speech Continuation Examples
@@
Appendix E Number of Iteration for Curriculum Learning
@@
Appendix F Instruction for the Mistral LLM for evaluation
@@
Appendix G Instruction for the GPT4-o for evaluation
@@
Appendix H auto-BLEU Score Distribution
@@
Appendix I Mistral Score Distribution
@@
Appendix J Correlation Between GPT4-o score and MMOS